# GLUE: Gradient-free Learning to Unify Experts

### v1

@@ Abstract
In many deployed systems (multilingual ASR, cross-hospital imaging, region-specific perception), multiple pretrained specialist models coexist. Yet, new target domains often require domain expansion: a generalized model that performs well beyond any single specialist’s domain. Given such a new target domain, prior works seek a single strong initialization prior for the model parameters by first blending expert models to initialize a target model. However, heuristic blending—using coefficients based on data size or proxy metrics—often yields lower target-domain test accuracy, and learning the coefficients on the target loss typically requires computationally-expensive full backpropagation through the network. We propose GLUE, Gradient-free Learning To Unify Experts, which initializes the target model as a convex combination of fixed experts, learning the mixture coefficients of this combination via a gradient-free two-point (SPSA) update that requires only two forward passes per step. Across experiments on three datasets and three network architectures, GLUE produces a single prior that can be fine-tuned effectively to outperform baselines. GLUE improves test accuracy by up to 8.5% over data-size weighting and by up to 9.1% over proxy-metric selection. GLUE either outperforms backpropagation-based full-gradient mixing or matches its performance within 1.4%.
@@ 1 Introduction
Pretrained parameters, whether from specialist “experts” or large foundation models, enable practitioners to start learning new tasks quickly by initializing a target network using the pretrained parameters and then finetuning it on the new task data [noroozi2018boosting, you2021logme, park2025fedbaf]. This practice is commonplace in multilingual or multi-accent automatic speech recognition (ASR) [turan2020achieving], radiology models deployed across scanners and hospitals [akbarian2023evaluating], perception stacks that face region/weather variation [shu2022hub, pal2024transfer], recommendation systems partitioned by locale or user segments [fu2024exploring], and foundation model adapters specialized to code, legal, or biomedical text [yuan2023power, zhou2024comprehensive, park2025fedbaf]. In these settings, initializing the target network with expert parameters often outperforms knowledge distillation [hinton2015distilling, chaudhari2023knowledge], which can introduce optimization noise (e.g., temperature tuning, teacher-student mismatch) and requires extra hyperparameters. In contrast, directly initializing the parameters of the target network with those of the experts simplifies downstream fine-tuning while retaining the pretrained feature hierarchy and inductive biases [zagoruyko2016paying, micaelli2019zero, mirzadeh2020improved, moslemi2024survey].
In general, adopting the parameters of a single expert is often suboptimal when building a general-purpose model, because real-world data are fragmented across related but distinct distributions [cohenask, shi2022robust, xiang2020learning]. Different experts capture complementary structures [yuan2021reinforced, wu2021one, boiarov2022simultaneous], motivating a parameter-space fusion of these experts as an initialization prior. Previous works [mcmahan2017communication, NonIID_P2P_Asilomar, 5GIoT_P2P_Asilomar, NTK_P2P_ICASSP, wortsman2022model, shi2025fedawa] hence seek to blend $K$ expert parameters $\{\boldsymbol{\theta}_{i}\}_{i=1}^{K}$ as
@@
Common heuristics for finding $\boldsymbol{\alpha}$ include (i) weighting experts by the size of their pretraining data [mcmahan2017communication], (ii) weighting by an external metric such as held-out accuracy on a proxy dataset[wortsman2022model, matena2022merging, park2025fedbaf], or (iii) learning the weights by optimizing the target loss[ilharcoediting, yadav2023ties, yangadamerging]. Among these, learning the weights $\boldsymbol{\alpha}$ generally leads to the strongest single initialization (prior) for fine-tuning, because it optimizes the true target loss and can capture interactions among experts that global heuristics miss [mu2025comprehensive, wortsman2022model].
The practical challenge, however, is cost. Even when the number of mixture weights $K$ is small, standard gradient descent on the mixture coefficients still requires computing the loss gradient $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$ for the full parameter vector $\boldsymbol{\theta}(\boldsymbol{\alpha})$ of dimension $P>>K$, thus requiring a complete forward and backward pass over the new model.
We propose Gradient-free Learning to Unify Experts (GLUE), a method for learning expert mixing coefficients $\boldsymbol{\alpha}$ via two-point simultaneous perturbation stochastic approximation (SPSA) [sadegh2002optimal, gu2021black]. Starting from (1), GLUE optimizes only the low-dimensional vector $\boldsymbol{\alpha}$ [jacobs1991adaptive, liudarts]. To do so, at each iteration, it samples a random direction $\mathbf{u}\in\mathbb{R}^{K}$, evaluates the loss at $\boldsymbol{\alpha}\pm\delta\mathbf{u}$ (two forward passes with blended weights $\boldsymbol{\theta}(\boldsymbol{\alpha}\pm\delta\mathbf{u})$), and updates $\boldsymbol{\alpha}$ using the finite-difference estimate. No backpropagation is required—only forward passes—making GLUE lightweight and easy to deploy. While SPSA has been used in other learning tasks, it may be inaccurate in high dimensions [sadegh2002optimal, nesterov2017random, gu2021black]. The key insight behind GLUE is that the mixture coefficients are only $K$-dimensional, mitigating this common drawback of SPSA. Our contributions are threefold:
@@ 2 Related Work
A straightforward way to combine experts is to weight each model in proportion to the amount of data used to train it—an idea resembling that of federated averaging (FedAvg), where model updates from multiple clients are aggregated based on clients’ local dataset sizes in order to produce a unified model that (on average) performs well on all clients’ data [mcmahan2017communication]. While simple and scalable, this strategy is agnostic to the target domain. Under distribution changes, such as domain expansion, data-size weighting cannot adapt the mixture to target loss signals and can overemphasize large but misaligned sources.
Another common approach involves choosing or reweighting experts using a proxy metric (e.g., held-out accuracy) and then averaging the expert model parameters. Examples include “model soups”—uniform averaging or greedy selection before averaging—and Fisher-weighted merging, which uses per-parameter Fisher information with model-level mixing as a hyperparameter  [wortsman2022model, matena2022merging, park2025fedbaf]. These methods can be effective when a reliable proxy set exists, but they decouple weight choice from the training objective and often require manual tuning or discrete selection steps.
A third line of work learns the expert mixture coefficients by minimizing a target-driven objective, ranging from scalar task-vector combinations to interference-aware merging and adaptive (un)supervised schemes that tune model-wise coefficients [ilharcoediting, yadav2023ties, yangadamerging]. However, these gradient-based methods require backpropagating through the blended network to compute $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$ and then taking inner products with the expert parameters (e.g., $\partial\mathcal{L}/\partial\boldsymbol{\alpha}_{i}=\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},\,\boldsymbol{\theta}_{i}\rangle$), which is computationally expensive.
Zeroth-order methods replace backpropagation with finite differences. In the two-point SPSA estimator [spall2002multivariate], we perturb $\boldsymbol{\alpha}$ along a random direction and estimate a directional derivative with two forward passes, staying in the $K$-dimensional mixture space where $K$ is the number of experts. However, the estimator’s variance grows with dimension and curvature, demanding many directions or delicate tuning and often causing slow or unstable convergence for large $K$ or ill-conditioned losses [ZHANG2022110006].
@@ 3.1 Objective
Let $f(\cdot;{\boldsymbol{\theta}})$ denote the network instantiated with parameters $\boldsymbol{\theta}$. For a minibatch $\mathcal{B}=\{(\mathbf{x}_{j},\mathbf{y}_{j})\}_{j=1}^{b}$ from the target domain and a standard supervised loss $\ell$ (cross-entropy for classification or squared error for regression), we minimize the empirical risk:
$\mathcal{L}(\boldsymbol{\alpha})=\frac{1}{b}\sum_{j=1}^{b}\ell\left(f(\mathbf{x}_{j};{\boldsymbol{\theta}(\boldsymbol{\alpha})}),\,\mathbf{y}_{j}\right)$ (2)
@@ 3.2 Two-point SPSA updates
Because our optimization variable is the $K$-dimensional mixture, and zeroth-order estimates are accurate in lower dimensions [sadegh2002optimal, nesterov2017random, gu2021black], we estimate the gradient of $\mathcal{L}$ with respect to $\boldsymbol{\alpha}$ using a two-point simultaneous perturbation.
At each iteration, we draw a random direction vector $\mathbf{u}\in\mathbb{R}^{K}$ (e.g., by drawing entries of $\mathbf{u}$ i.i.d. from a Gaussian or Rademacher distribution and normalizing) and choose a small radius $\mu>0$. We form perturbed unconstrained coefficients $\boldsymbol{\alpha}_{\pm}=\boldsymbol{\alpha}\pm\mu\mathbf{u}$ and corresponding blended parameters $\boldsymbol{\theta}(\boldsymbol{\alpha}_{\pm})=\sum_{i=1}^{K}{\alpha}_{{\pm}_{i}}\,\boldsymbol{\theta}_{i}$. At both blends, we then evaluate the loss on the same minibatch of target-domain examples, keeping inference deterministic across the two probes (e.g., BatchNorm in evaluation mode and fixed randomness) so that differences reflect only the perturbation:
@@
$d(\boldsymbol{\alpha};\mathbf{u})=\frac{L_{+}-L_{-}}{2\mu}$ (4)
@@
$\widehat{\nabla}_{\boldsymbol{\alpha}}\mathcal{L}(\boldsymbol{\alpha})=d(\boldsymbol{\alpha};\mathbf{u})\,\mathbf{u}$ (5)
When desired, we can reduce variance by additionally averaging over $m$ independently sampled directions $\{\mathbf{u}_{r}\}_{r=1}^{m}$ in the same iteration,
@@
where $\eta>0$ is the step size. In our experiments, we perform SPSA and momentum updates in the unconstrained $\boldsymbol{\beta}$-space and map to $\boldsymbol{\alpha}=\mathrm{softmax}(\boldsymbol{\beta})$, replacing (7). Optimizing in $\boldsymbol{\beta}$ ensures that $\boldsymbol{\theta}(\boldsymbol{\alpha})$ remains within the convex hull of the expert parameters.
@@ 3.3 From mixture to prior
We initialize $\boldsymbol{\alpha}=\frac{1}{K}\mathbf{1}$ and iterate the two-point updates over minibatches until validation performance plateaus. Formally, we target the empirical-risk minimizer and then map it to a single blended model:
$\boldsymbol{\alpha}^{\star}\in\operatorname*{argmin}_{\boldsymbol{\alpha}\in\mathbb{R}^{K}}\mathcal{L}(\boldsymbol{\alpha}),\;\qquad\boldsymbol{\theta}^{\star}=\boldsymbol{\theta}(\boldsymbol{\alpha}^{\star})$ (8)
The resulting $\boldsymbol{\theta}^{\star}$ serves as the starting point for standard gradient-based adaptation on the target domain, providing a strong prior that concentrates the most relevant inductive biases from the experts.
@@ 4 Theoretical Analysis
In this section, we show that the two-point (SPSA) procedure used in GLUE is (i) computationally cheaper per iteration than gradient-based mixing and (ii) statistically stable when the mixture dimension $K$ (the number of experts) is sufficiently small. To do so, we leverage results in stochastic approximation and zeroth-order optimization [sadegh2002optimal, nesterov2017random, gu2021black].
4.1 Two-point (SPSA) vs. full-gradient cost for updating $\boldsymbol{\alpha}$}
@@
Let $F$ denote the cost of one forward pass of the blended network on the current minibatch, and let the cost of a full backward pass be $\gamma F$ with $\gamma\geq 2$ (typical for deep neural networks [gruslys2016memory]). Let $C_{\text{mix}}$ be the cost of the multiply-accumulate operations used to form a blended parameter vector ${\boldsymbol{\theta}}({\boldsymbol{\alpha}})$ from the $K$ experts: $C_{\text{mix}}=O(PK)$ memory operations for parameter dimension $P$. Let $D_{{\boldsymbol{\alpha}}}$ be the cost of computing the gradient $\nabla_{\boldsymbol{\alpha}}\mathcal{L}=\big[\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},{\boldsymbol{\theta}}_{1}\rangle,\ldots,\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},{\boldsymbol{\theta}}_{K}\rangle\big]$ given $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$, which is an $O(PK)$ dot-product accumulation. Evaluating $\nabla_{\boldsymbol{\alpha}}\mathcal{L}$ at the current mixture requires (i) computing ${\boldsymbol{\theta}}({\boldsymbol{\alpha}})$ once, (ii) one forward pass and one backward pass to compute $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$, and (iii) $D_{\boldsymbol{\alpha}}$ for computing $K$ inner products with $\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},\boldsymbol{\theta}_{i}\rangle$. With one perturbation pair ($m=1$), SPSA evaluates the loss at two nearby mixtures ${\boldsymbol{\alpha}}_{\pm}={\boldsymbol{\alpha}}\pm\delta\mathbf{u}$, requiring two blends and two forwards, but no backward. Therefore:
$T_{\text{full}}$ $=(1+\gamma)F+C_{\text{mix}}+D_{{\boldsymbol{\alpha}}},\qquad T_{\text{SPSA}}=2(F+C_{\text{mix}})$ (9)
Whenever $T_{\text{SPSA}}<T_{\text{full}}$, SPSA is cheaper than the full-gradient step. Therefore, whenever
$C_{\text{mix}}$ $<(\gamma-1)F+D_{\boldsymbol{\alpha}}$ (10)
a single SPSA pair is cheaper than a full-gradient step. In practice, the condition is easily met since (i) the inner product cost $D_{\boldsymbol{\alpha}}$ is typically of the same order as the mixing cost $C_{\text{mix}}$ and (ii) the cost of the backward pass is typically more than $2F$ ($\gamma\geq 2$).

### v2

@@ Abstract
In many deployed systems (multilingual ASR, cross-hospital imaging, region-specific perception), multiple pretrained specialist models coexist. Yet, new target domains often require domain expansion: a generalized model that performs well beyond any single specialist’s domain. Given a new target domain, existing methods obtain a single strong initialization prior for the model parameters by blending expert models to initialize a target model. However, heuristic blending—using mixing coefficients based on data size or proxy metrics—often yields lower target-domain test accuracy, and learning these coefficients on the target domain’s loss function typically requires computationally-expensive full backpropagation through a neural network. We propose GLUE, Gradient-free Learning to Unify Experts, which initializes the target model as a convex combination of fixed experts and learns the mixture coefficients of this combination via gradient-free two-point SPSA (simultaneous perturbation stochastic approximation) updates, requiring only two forward passes per step. Across experiments on three datasets and three network architectures, GLUE produces model parameter priors that can be fine-tuned to outperform baselines. GLUE improves test accuracy by up to 8.5% over data-size weighting and by up to 9.1% over proxy-metric selection. GLUE either outperforms backpropagation-based full-gradient mixing or matches its performance within 1.4%.
@@ 1 Introduction
Pretrained parameters, whether from specialist “expert” models or large foundation models, enable practitioners to quickly start learning new tasks by using the pretrained parameters to initialize a target network and finetune it on the new task data [22, 38, 24]. This practice is common in multilingual or multi-accent automatic speech recognition (ASR) [32], radiology models deployed across scanners and hospitals [1], perception stacks that face region/weather variation [30, 23], recommendation systems partitioned by locale or user segments [7], and foundation model adapters specialized for code, legal, or biomedical text [40, 43, 24]. In these settings, initializing a target network with expert parameters often outperforms knowledge distillation [11, 4], which requires extra hyperparameters and can introduce optimization noise (e.g., temperature tuning, teacher-student mismatch). In contrast, directly initializing a target network with experts’ parameters simplifies downstream fine-tuning while preserving pretrained feature hierarchy and inductive biases [41, 17, 18, 19].
In general, adopting a single expert’s parameters is often suboptimal when building a general-purpose model, because real-world data are fragmented across related but distinct distributions [6, 29, 35]. Different experts capture complementary structures [39, 34, 3], motivating a parameter-space fusion of these experts as an initialization prior. Previous works [16, 25, 26, 5, 33, 28] form convex combinations of the $P$-dimensional parameters of $K$ expert models $\{\boldsymbol{\theta}_{i}\}_{i=1}^{K}$
@@
Common heuristics for finding $\boldsymbol{\alpha}$ include (i) weighting experts by the size of their pretraining data [16], (ii) weighting by an external metric such as held-out accuracy on a proxy dataset[33, 15, 24], or (iii) learning the weights by optimizing the target loss[12, 36, 37]. Among these, learning the weights $\boldsymbol{\alpha}$ generally leads to the strongest single initialization (prior) for fine-tuning, because it optimizes the true target loss and can capture interactions among experts that global heuristics miss [20, 33].
The practical challenge is computational cost. Even when the number of mixture weights $K$ is small, standard gradient descent on the mixture coefficients still requires computing the loss gradient $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$ for the full parameter vector $\boldsymbol{\theta}(\boldsymbol{\alpha})$ of dimension $P\gg K$, thus requiring a complete forward and backward pass over the new model.
We propose Gradient-free Learning to Unify Experts (GLUE), a method for learning expert mixing coefficients $\boldsymbol{\alpha}$ via two-point simultaneous perturbation stochastic approximation (SPSA) [27, 10]. Starting from (1), GLUE optimizes only the low-dimensional vector $\boldsymbol{\alpha}$ [13, 14]. To do so, at each iteration, it samples a random direction $\mathbf{u}\in\mathbb{R}^{K}$, evaluates the loss at $\boldsymbol{\alpha}\pm\delta\mathbf{u}$ (two forward passes with blended weights $\boldsymbol{\theta}(\boldsymbol{\alpha}\pm\delta\mathbf{u})$), and updates $\boldsymbol{\alpha}$ using the finite-difference estimate. No backpropagation is required—only forward passes—making GLUE lightweight and easy to deploy. While SPSA has been used in other learning tasks, it may be inaccurate in high dimensions [27, 21, 10]. The key insight behind GLUE is that the mixture coefficients are only $K$-dimensional, mitigating this common drawback of SPSA. Our contributions are threefold:
@@ 2 Related Work
A straightforward way to combine experts is to weight each model in proportion to the amount of data used to train it—an idea resembling that of federated averaging (FedAvg), where model updates from multiple clients are aggregated based on clients’ local dataset sizes in order to produce a unified model that (on average) performs well on all clients’ data [16]. While simple and scalable, this strategy is agnostic to the target domain. Under distribution changes, such as domain expansion, data-size weighting cannot adapt to target loss signals and can overemphasize large but misaligned sources.
Another common approach involves choosing or reweighting experts using a proxy metric (e.g., held-out accuracy) and then averaging the expert model parameters. Examples include “model soups”—uniform averaging or greedy selection before averaging—and Fisher-weighted merging, which uses per-parameter Fisher information with model-level mixing as a hyperparameter  [33, 15, 24]. These methods can be effective when a reliable proxy set exists, but they decouple weight choice from the training objective and often require manual tuning or discrete selection steps.
A third line of work learns the expert mixture coefficients by minimizing a target-driven objective, ranging from scalar task-vector combinations (TIES-Merging [36]) to interference-aware merging and adaptive (un)supervised schemes that tune model-wise coefficients (AdaMerging) [12, 37]. These gradient-based methods require computing $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$ via backpropagation through the blended network and then taking inner products with expert parameters (e.g., $\partial\mathcal{L}/\partial\boldsymbol{\alpha}_{i}=\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},\,\boldsymbol{\theta}_{i}\rangle$), which is computationally expensive.
Zeroth-order methods replace backpropagation with finite differences. To compute the two-point simultaneous perturbation stochastic approximation (SPSA) estimator [31], $\boldsymbol{\alpha}$ can be perturbed along a random direction to estimate a directional derivative with two forward passes, staying in the $K$-dimensional mixture space (where $K$ is the number of experts). However, the estimator’s variance grows with dimension and curvature, demanding many directions or delicate tuning and often causing slow or unstable convergence for large $K$ or ill-conditioned losses [42].
@@ 3.1 Objective
Let $f(\cdot;{\boldsymbol{\theta}})$ denote the model (neural network) instantiated with parameters $\boldsymbol{\theta}$. For a minibatch $\mathcal{B}=\{(\mathbf{x}_{j},\mathbf{y}_{j})\}_{j=1}^{B}$ from the target domain and a standard supervised loss $\ell$ (cross-entropy for classification or squared error for regression), we minimize empirical risk:
$\mathcal{L}(\boldsymbol{\alpha})=\frac{1}{B}\sum_{j=1}^{B}\ell\left(f(\mathbf{x}_{j};{\boldsymbol{\theta}(\boldsymbol{\alpha})}),\,\mathbf{y}_{j}\right)$
@@ 3.2 Two-point SPSA updates
Since our optimization variable is the $K$-dimensional mixture, and zeroth-order estimates are accurate in lower dimensions [27, 21, 10], we estimate the gradient of $\mathcal{L}$ with respect to $\boldsymbol{\alpha}$ by using a two-point simultaneous perturbation stochastic approximator (SPSA).
At each iteration, we draw a random direction vector $\mathbf{u}\in\mathbb{R}^{K}$ (e.g., by drawing entries of $\mathbf{u}$ i.i.d. from a standard Gaussian distribution and normalizing) and choose a small radius $\mu>0$. We form perturbed unconstrained coefficients $\boldsymbol{\alpha}_{\pm}=\boldsymbol{\alpha}\pm\mu\mathbf{u}$ and corresponding blended parameters $\boldsymbol{\theta}(\boldsymbol{\alpha}_{\pm})=\sum_{i=1}^{K}{\alpha}_{{\pm}_{i}}\,\boldsymbol{\theta}_{i}$. At both blends, we then evaluate the loss on the same minibatch of target-domain examples, keeping inference deterministic across the two probes (e.g., BatchNorm in evaluation mode and fixed randomness) so that differences reflect only the perturbation:
@@
$d(\boldsymbol{\alpha};\mathbf{u})=\frac{L_{+}-L_{-}}{2\mu}$
@@
$\widehat{\nabla}_{\boldsymbol{\alpha}}\mathcal{L}(\boldsymbol{\alpha})=d(\boldsymbol{\alpha};\mathbf{u})\,\mathbf{u}$
If desired, we can reduce variance by additionally averaging over $m$ independently sampled directions $\{\mathbf{u}_{r}\}_{r=1}^{m}$ in the same iteration,
@@
where $\eta>0$ is the step size. In our experiments, we perform SPSA and momentum updates in an unconstrained $\boldsymbol{\beta}$-space and map to $\boldsymbol{\alpha}=\mathrm{softmax}(\boldsymbol{\beta})$, replacing (4). Optimizing with respect to $\boldsymbol{\beta}$ ensures that $\boldsymbol{\theta}(\boldsymbol{\alpha})$ remains within the convex hull of expert parameters.
@@ 3.3 From mixture to prior
We initialize $\boldsymbol{\alpha}=\frac{1}{K}\mathbf{1}$ and iterate the two-point updates over minibatches until validation performance plateaus. Formally, we target the empirical-risk minimizer and map it to one blended model:
$\boldsymbol{\alpha}^{\star}\in\operatorname*{argmin}_{\boldsymbol{\alpha}\in\mathbb{R}^{K}}\mathcal{L}(\boldsymbol{\alpha}),\;\qquad\boldsymbol{\theta}^{\star}=\boldsymbol{\theta}(\boldsymbol{\alpha}^{\star})$
The resulting $\boldsymbol{\theta}^{\star}$ serves as the starting point for standard gradient-based adaptation on the target domain, providing a strong prior that concentrates the experts’ most relevant inductive biases.
@@ 4 Theoretical Analysis
In this section, we show that the two-point (SPSA) procedure used in GLUE is (i) computationally cheaper per iteration than gradient-based mixing and (ii) statistically stable when the mixture dimension $K$ (the number of experts) is sufficiently small. To do so, we leverage results in stochastic approximation and zeroth-order optimization [27, 21, 10]; under standard smoothness and bounded-noise assumptions, two-point SPSA converges to first-order stationary points, so GLUE inherits these classical guarantees [8].
4.1 Cost for updating $\boldsymbol{\alpha}$}: two-point (SPSA) vs. full-gradient
@@
Let $F$ denote the cost of one forward pass of the blended network on a minibatch of data, and let the cost of a full backward pass be $\gamma F$ with $\gamma\geq 2$ (typical for deep neural networks [9]). Let $C_{\text{mix}}$ be the cost of the multiply-accumulate operations used to form a blended parameter vector ${\boldsymbol{\theta}}({\boldsymbol{\alpha}})$ from the $K$ experts: $C_{\text{mix}}=O(PK)$ memory operations for parameter dimension $P$. Let $D_{{\boldsymbol{\alpha}}}$ be the cost of computing the gradient $\nabla_{\boldsymbol{\alpha}}\mathcal{L}=\big[\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},{\boldsymbol{\theta}}_{1}\rangle,\ldots,\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},{\boldsymbol{\theta}}_{K}\rangle\big]$ given $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$, which is an $O(PK)$ dot-product accumulation. Evaluating $\nabla_{\boldsymbol{\alpha}}\mathcal{L}$ at the current mixture requires (i) computing ${\boldsymbol{\theta}}({\boldsymbol{\alpha}})$ once, (ii) one forward pass and one backward pass to compute $\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L}$, and (iii) $D_{\boldsymbol{\alpha}}$ for computing $K$ inner products with $\langle\nabla_{\boldsymbol{\theta}(\boldsymbol{\alpha})}\mathcal{L},\boldsymbol{\theta}_{i}\rangle$. With one perturbation pair ($m=1$), SPSA evaluates the loss at two nearby mixtures ${\boldsymbol{\alpha}}_{\pm}={\boldsymbol{\alpha}}\pm\delta\mathbf{u}$, requiring two blends and two forwards, but no backward. Therefore:
$T_{\text{full}}$ $=(1+\gamma)F+C_{\text{mix}}+D_{{\boldsymbol{\alpha}}},\qquad T_{\text{SPSA}}=2(F+C_{\text{mix}})$
Whenever $T_{\text{SPSA}}<T_{\text{full}}$, SPSA is cheaper than the full-gradient step (i.e., $C_{\text{mix}}<(\gamma-1)F+D_{\boldsymbol{\alpha}}$). In practice, the condition is easily met since (i) the inner product cost $D_{\boldsymbol{\alpha}}$ is typically of the same order as the mixing cost $C_{\text{mix}}$ and (ii) the cost of the backward pass is typically more than $2F$ ($\gamma\geq 2$).